OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. You shoot a model rocket into the air with a speed of 18.2 m/s. How high does the rocket go? A. 16.9 m B. 0.93 m C. 8.2 m D. 4.5 m

OpenStudy (anonymous):

Ok, that's simple. You are given the initial velocity v1 = 18.2 m/s, the acceleration due to gravity a = -9.8 m/s. Since the rocket is launching upward against gravity, the value of g will be negative. From your question, it looks like you want to find the max. Height the rocket reaches. At max. Height for a moment the final velocity v2 will be 0 m/s. V1 = 18.2 m/s a= -9.8 m/s^2 V2 = 0 m/s Since we don't have the time, use one of the big five kinematic equations to solve for the height by using the given values. Good luck! ( in the first paragraph, I meant to say a = -9.8 m/s^2)

OpenStudy (anonymous):

so, what equation would I use?

OpenStudy (anonymous):

@coolsday

OpenStudy (anonymous):

V2^2 = v1^2 + 2ad

OpenStudy (anonymous):

Isolate d from the formula to find the max. Height

OpenStudy (anonymous):

The five equations should be in ur notes or textbook

OpenStudy (anonymous):

so, is the answer A? @coolsday

OpenStudy (anonymous):

Correct